a model cylinder fuel tank has a radius of 5cm and a height of 12cm. Scale is 1cm to 1.25m.
How tall is the actual fuel tank in meters? 15m?
What is the scale factor between model and fuel tank?1/150?
Determine the volume in cubic meters. Determine surface area of the fuel tank in sq m.
I don't mind paYING A LITTLE MONEY, i JUST DON'T HAVE ALOT. PLEASE HELP AN OL' LADY WHO IS TRYING TO GRADUATE HIGH SCHOOL AT 56. PLKEASE LET ME KNOW HOW MUCH AND WE CAN WORK SOMETHING OUT. gOD BLESS.

Found 2 solutions by jim_thompson5910, jerryguo41:
Answer by jim_thompson5910(28595)

1 cm : 1.25 m

5*1 cm : 5*1.25 m

5 cm : 6.25 m

----------------------------------------
1 cm : 1.25 m

12*1 cm : 12*1.25 m

12 cm : 15 m

----------------------------------------

So the actual tank has a radius of 6.25 meters and is 15 meters tall

Now convert 15 meters to cm: 15*100 = 1500 cm

Then divide it by 12 cm: 1500/12 = 125

So the scale factor is 1/125

-------------------------------------------------------

Volume:

V = pi*r^2*h

V = pi*(6.25)^2*15

V = pi*39.0625*15

V = pi*585.9375

V = 585.9375pi

V = 585.9375(3.14159)

V = 1840.775390625

So the volume of the tank is approximately 1840.775390625 cubic meters

Surface area

SA = 2*pi*r^2 + 2pi*r*h

SA = 2*pi*(6.25)^2 + 2pi*6.25*15

SA = 2*pi*39.0625 + 2pi*6.25*15

SA = 78.125pi + 187.5pi

SA = 265.625pi

SA = 265.625(3.14159)

SA = 834.48484375

and the surface area is approximately 834.48484375 square meters.
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------

Answer by jerryguo41(162)

How tall is the actual fuel tank in meters?
Model: 12cm
12cm*(1.25m/1cm)=15m
Actual:15m
---------------------------------------------------------
What is the scale factor between model and fuel tank?
Model: 1cm
Actual: 1.25m or 125 cm
Ration: model/actual -> 1/125
---------------------------------------------------------
Determine the volume in cubic meters.
Find Radius of Actual:
Model: 5cm
5cm*(1.25m/1cm)=6.25m
Actual:6.25m
---------------------------------------------------------
Volume of Cylinder:V=πr^2*h (r=radius)(h=height)(V=Volume)
V=π6.25^2*15
V=π39.0625*15
V=585.9375π or Approx. 1,839.84 cubic meters
---------------------------------------------------------
Surface Area of a Cylinder: Sa=2πr^2+2πr*h (r=radius)(h=height)(Sa=Surface Area)
Sa=2π6.25^2+2π6.25*15
Sa=2π39.0625+2π93.75
Sa=78.125π+187.5π
Sa=265.625π or 834.0625 Square meters